What mass of liquid water can be heated from freezing to boiling with 42 kJ of energy, if the specific heat of water is 4.2 J g-1 oC-1?

1
Expert's answer
2021-04-21T03:54:09-0400

H= mcΔt

m= H/cΔt

= 42000/(4.2×100) = 100g